# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ 3.1.1 Bulk convection speeds
Figure 4: (a) The scatter plot between the convection speed estimated from the DTS measurements ($U_{C}$) and the mean wind speed from EC system ($U$) at the clearcut. The dash-dotted black line show the 1:1 relationship while the blue dash-dotted line is the best fit line. (b)–(c) The ratios $U_{c}/U$ are plotted against the turbulence intensities $\sigma_{u}/U$ for the cases when the direction of the mean wind was either along or perpendicular to the forest edge. For the perpendicular case, the wind directions are further subdivided when the wind was approaching directly from the forest or from the clearcut. (d) The ratios between the convection speeds of positive ($U^{+}_{c}$) and negative ($U^{-}_{c}$) temperature structures are shown against the skewness of temperature fluctuations obtained from the EC measurements. (e) By combining the EC and DTS measurements, a scatter plot between the length scales of positive and negative temperature structures is shown. Here, $z$ is the measurement height, $\gamma^{+}_{T}$ and $\gamma^{-}_{T}$ are the integral time scales of positive and negative temperature fluctuations while $U^{+}_{c}$ and $U^{-}_{c}$ denote their convection speeds, respectively.
@@
According to the study by Willis and Deardorff (1976), the Taylor’s hypothesis is rendered invalid when the turbulent intensities of the streamwise velocities ($\sigma_{u}/U$, where $\sigma_{u}$ is the standard deviation) exceed 0.5. In Figs. 4b–c, we show the scatter plots between $U_{c}/U$ and $\sigma_{u}/U$. Although the number of samples are limited, it appears that for the along-wind cases (Fig. 4b), $U_{c}/U$ indeed remains close to 1 when $\sigma_{u}/U\leq 0.5$. However, for values other than 0.5, the relationship between $U_{c}/U$ and $\sigma_{u}/U$ is linear with an asymptote of 1.9 as $\sigma_{u}/U>0.8$. All these relationships can be condensed to a singular form,
$U_{c}/U=\mathcal{H}(0.5-\sigma_{u}/U)+(3\sigma_{u}/U-0.5)[\mathcal{H}(\sigma_{u}/U-0.5)-\mathcal{H}(\sigma_{u}/U-0.8)]+1.9\mathcal{H}(\sigma_{u}/U-0.8),$ (7)
@@
For the perpendicular-wind cases, the mean wind directions can be further separated into two categories; when the wind was blowing directly from the forest and when it approached the forest edge from the clearcut. For the first category (forest), it appears that the ratios $U_{c}/U$ were closer to unity despite the changes in $\sigma_{u}/U$ (Fig. 4c). Therefore, corresponding to this flow category, the Taylor’s hypothesis appears to be valid even for $\sigma_{u}/U>0.5$. However, for the latter category (clearcut), the pink squares in Fig. 4c followed Eq. 7 more closely. This difference is not trivial by considering the fact that when the wind blows from the forest, the turbulence at the clearcut location has not fully equilibrated to the local surface conditions. On the other hand, when the wind approaches the forest edge from the clearcut, the situation reverses.
[…]
3.1.2 Conditional convection speeds
Moving on, to see whether the positive and negative temperature structures move at the same speeds, we computed the ratio $U^{+}_{C}/U^{-}_{C}$, where $U^{+}_{C}$ and $U^{-}_{C}$ were estimated from the conditional space-time correlation curves (Figs. 3b–d). From a statistical perspective, the asymmetry between positive and negative temperatures is related to the skewness of temperature fluctuations. As shown in Fig. 9a of Appendix A, for our selected periods, the distributions of temperature fluctuations are positively skewed. In Fig. 4d, the ratios $U^{+}_{C}/U^{-}_{C}$ were plotted against the skewness of temperature fluctuations ($\mathcal{S}(T)$) computed from the EC measurements. This was done to check if the statistical asymmetry could explain any differences between $U^{+}_{C}$ and $U^{-}_{C}$. Note that in Fig. 4d there are fewer data-points corresponding to the perpendicular-wind cases. This is because the peaks ($\tau^{\pm}_{p}$) of the conditional space-time correlation curves did not move nicely with $r$ in order to have robust estimates of $U^{+}_{C}$ and $U^{-}_{C}$. For the along-wind cases, $U^{+}_{C}/U^{-}_{C}$ values were averaged over all the five sections and the spread between them is shown as error-bars with one standard deviation. From Fig. 4d it appears that, for both along- and perpendicular-wind cases, the ratios $U^{+}_{C}/U^{-}_{C}$ remain at a constant value of 0.8 irrespective of temperature skewness.
Since $U^{+}_{C}/U^{-}_{C}<1$, the up-plumes tend to move slower than the down-plumes. In essence, this result suggests that the temporal statistics are based on a biased sample of turbulent flow when compared against sampling over spatial domain, which, however, is not in practical terms possible. In terms of boundary-layer physics, the down-plumes bring colder air from aloft to the layer close to the surface and since these plumes carry the signatures of the large-scale structures, they appear to move faster (Li et al., 2018). We now use this critical information to compute the averaged sizes of up- and down-plumes from the EC measurements. To do that, we first estimate the auto-correlation functions conditioned upon whether the temperatures are positive or negative. Mathematically, these functions can be defined as,
$R^{+}_{TT}(\tau)=\frac{\overline{[T^{\prime}(t)|>0][T^{\prime}(t+\tau)|>0]}}{\sigma^{2}_{T}},$ (8)
where $R^{+}_{TT}(\tau)$ indicate the autocorrelation functions of temperature fluctuations conditioned on their positive values only. $R^{-}_{TT}(\tau)$ is defined in a similar way but for the negative ones. By integrating $R^{+}_{TT}(\tau)$ and $R^{-}_{TT}(\tau)$ up to their first zero crossing, one can estimate the integral time scales of the positive ($\gamma^{+}_{T}$) and negative ($\gamma^{-}_{T}$) temperature structures. The functions $R^{+}_{TT}(\tau)$ and $R^{-}_{TT}(\tau)$ are compared with the original autocorrelation functions ($R_{TT}(\tau)$) in Fig. S3 of the Supplementary Material. By multiplying $\gamma^{+}_{T}$ and $\gamma^{-}_{T}$ with their respective convection speeds ($U^{+}_{C}$ and $U^{-}_{C}$), the physical dimensions of the up- and down-plumes can be calculated.
In Fig. 4e, we show a scatter plot between $(\gamma^{+}_{T}U^{+}_{C})/z$ and $(\gamma^{-}_{T}U^{-}_{C})/z$, where $z$ is the measurement height ($z=3.1$ m). We only show this comparison for the along-wind cases and one can clearly see that the sizes of the down-plumes can be as large as 12 times the measurement height. In a statistical sense, the down-plumes remain nearly 2.5 times larger than the up-plumes (see the blue dash-dotted line in Fig. 4e). This information is important e.g. from the perspective of choosing the domain size for the large eddy simulation models, since the down-plumes would require considerably larger area for their statistics to converge.
@@ 3.1.3 Scale-dependent convection speeds
Figure 5: (a) The scale-dependent convection speeds ($U_{c}(\kappa z)$) are plotted against the streamwise wavenumbers ($\kappa$) computed directly from the DTS spatial information. The wavenumbers are normalized by the measurement height ($\kappa z$) while $U_{c}(\kappa z)$ values are scaled by the bulk convection speed $U_{C}$. The gray- and red-colored lines denote the along- and perpendicular-wind cases, respectively. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. The various fits to the scale-dependent convection speeds are shown by the dash-dotted lines (see the legend). (b) Same as in (a), but $U_{c}(\kappa z)$ values are scaled by the convection speed estimated from an elliptical model of space-time correlation curves. (c)–(d) The scale-dependent convection speeds are shown separately for the positive and negative temperature structures.
@@
Hitherto, we have discussed the bulk convective speeds but now we turn our attention towards the scale-dependent convection speeds. In Fig. 5a, we show the ensemble-averaged scale-dependent convection speed estimations from the spatial DTS data of temperature at all the six sections. As before, the Sections 1–5 correspond to along-wind cases while Section 6 represents the perpendicular ones. The convection speeds in Fig. 5a are normalized by the bulk convective speed $U_{C}$ and the wavenumbers ($\kappa$) computed from the spatial DTS separations are scaled with the measurement height ($z$). The choice of $z$ to scale $\kappa$ is in line with Hilland and Christen (2024), who claimed that in the context of Taylor’s hypothesis the measurement height is indeed an important parameter. Ideally, one should use $z-d$, where $d$ is the displacement height, but here we used only $z$ since to estimate $d$ a detailed vertical profile of wind measurements is required and that was not available at our measurement site. This would not have any impact on the results as $d$ only acts to shift the $x$ coordinates horizontally.
@@
Figure 5a depicts that at lower wavenumbers, $\kappa z<1$, $U_{C}(\kappa z)$ values tend towards $U_{C}$, irrespective of the wind directions. On the other hand, at larger wavenumbers, $\kappa z>1$, $U_{C}(\kappa z)/U_{C}$ values decrease in a power-law fashion with the exponent being equal to ${(\kappa z)}^{-0.8}$ (Fig. 5a). This is surprisingly different from the result of Cheng et al. (2017) who reported this exponent to be ${(\kappa z)}^{-1/3}$. We replotted the results of Cheng et al. (2017) and indeed for different atmospheric stability conditions (near-neutral to highly-convective) their data agree more closely with ${(\kappa z)}^{-1/3}$ instead of ${(\kappa z)}^{-0.8}$ (see Fig. S4b in the Supplementary material). However, compared to ours, the data of Cheng et al. (2017) exhibited a significant amount of scatter.
Cheng et al. (2017) theoretically derived this $-1/3$ exponent for atmospheric surface layer flows by exploiting the advection-diffusion equation and a power-law dependence of the diffusion coefficient under the assumption of homogeneous and isotropic turbulence at the inertial subrange scales. For our case, the apparent disagreement with their results could be attributed to two factors. First, we consider highly-heterogeneous flow conditions such as a forest clearcut. Second, the recent results of Chowdhuri and Banerjee (2024) have demonstrated that, for a wide range of both atmospheric and laboratory datasets, significant anisotropy can persist even at smaller scales of the flow. This is true for our EC dataset, as the flow remains anisotropic at the inertial subrange scales when computed through the intermittency-anisotropy framework of Chowdhuri and Banerjee (2024) (not shown).
For atmospheric surface layer and sparse canopy flows, a couple of previous studies have shown that an elliptic model of space-time correlation curves to compute the convection speeds is a better approach as opposed to a linear one adopted here (Han and Zhang, 2022; Everard et al., 2021). We describe this elliptical model in Appendix B. In this model, the convective speeds of the turbulent structures are estimated as $\sqrt{U_{e}^{2}+V_{e}^{2}}$, where $U_{e}$ and $V_{e}$ are convection and sweeping velocities, respectively. In Fig. 5b, $\sqrt{U_{e}^{2}+V_{e}^{2}}$ is used to normalize $U_{C}(\kappa z)$ and the results demonstrate that under this scaling the scale-dependent convection speeds do not exactly approach the bulk convective speeds ($\sqrt{U_{e}^{2}+V_{e}^{2}}$) at lower wavenumbers. Furthermore, the collapse among different sections is poorer as compared to Fig. 5a. Therefore, this questions the applicability of an elliptic model in these heterogeneous flow conditions.
To understand the origin of $-0.8$ exponent in scale-dependent convection speeds, we conditioned them based on the instances when the spatial temperature fluctuations were either positive or negative (see Sect. 2.5). For consistency, both $U^{+}_{C}(\kappa z)$ and $U^{-}_{C}(\kappa z)$ values were scaled with $U_{c}$. From Figs. 5c–d, it is clear that the $-0.8$ exponent is connected to the convective speeds of the down plumes ($U^{-}_{C}(\kappa z)$), contrary to the up ones ($U^{+}_{C}(\kappa z)$). In fact, at lower wavenumbers ($\kappa z<1$), $U^{-}_{C}(\kappa z)$ remain larger than $U^{+}_{C}(\kappa z)$ values and at larger wavenumbers ($\kappa z>1$) the convective speeds of the up plumes decrease as fast as ${(\kappa z)}^{-1.5}$. Therefore, as per the terminology of Cheng et al. (2017), the down-plumes diffuse slower than the up-plumes while carried along with the flow. We next discuss the implications of the scale-dependent convection speeds on the estimations of turbulence spectra and cospectra from EC measurements conducted in a temporal domain.
3.2 Impact on turbulence spectra

### v2

@@ 3.1.1 Bulk convection speeds
Figure 4: (a) The scatter plot between the convection speed estimated from the DTS measurements ($U_{C}$) and the mean wind speed from EC system ($U$) at the clearcut. The dash-dotted black line show the 1:1 relationship while the blue dash-dotted line is the best fit line. (b)–(c) The ratios $U_{C}/U$ are plotted against the turbulence intensities $\sigma_{u}/U$ for the cases when the direction of the mean wind was either along or perpendicular to the forest edge. For the perpendicular case, the wind directions are further subdivided when the wind was approaching directly from the forest or from the clearcut. (d) The scatter plot between $U_{e}$ and $U$ is shown, where $U_{e}$ is obtained from the $r_{p}$-$\tau$ relationship. The black dash dotted line indicates the 1:1 relationship. (e) The relationship between the square of the sweeping speed ($V^{2}$) and turbulence kinetic energy (TKE) is shown for the DTS cable sections 1–5 (along-wind conditions). The TKE is obtained from the EC system in the middle of the clearcut. The black, magenta, and red dash-dotted lines indicate the relationships $V^{2}\approx\rm TKE$, $V^{2}\approx\rm 2.16\ TKE$, and $V^{2}\approx\rm 2.25\ TKE$, respectively. The latter two are obtained from everard2021sweeping for flows over a vineyard canopy.
@@
According to the study by willis1976use, the Taylor’s hypothesis is rendered invalid when the turbulent intensities of the streamwise velocities ($\sigma_{u}/U$, where $\sigma_{u}$ is the standard deviation) exceed 0.5. In Figs. 4b–c, we show the scatter plots between $U_{C}/U$ and $\sigma_{u}/U$. Due to the scarcity in the number of samples, we cannot verify whether $U_{C}/U$ remains close to 1 when $\sigma_{u}/U\leq 0.5$. However, for values other than 0.5, the relationship between $U_{C}/U$ and $\sigma_{u}/U$ is linear with an asymptote of 1.9 as $\sigma_{u}/U>0.8$. All these relationships can be condensed to a singular form,
$U_{C}/U=(3\sigma_{u}/U-0.5)[\mathcal{H}(\sigma_{u}/U-0.5)-\mathcal{H}(\sigma_{u}/U-0.8)]+1.9\mathcal{H}(\sigma_{u}/U-0.8),$ (6)
@@
For the perpendicular-wind cases, the mean wind directions can be further separated into two categories; when the wind was blowing directly from the forest and when it approached the forest edge from the clearcut. For the first category (forest), it appears that the ratios $U_{C}/U$ were closer to unity despite the changes in $\sigma_{u}/U$ (Fig. 4c). However, for the latter category (clearcut), the pink squares in Fig. 4c followed Eq. 6 more closely. This difference is not trivial by considering the fact that when the wind blows from the forest, the turbulence at the clearcut location has not fully equilibrated to the local surface conditions. On the other hand, when the wind approaches the forest edge from the clearcut, the situation reverses.
To verify this argument, in Fig. S3 of the Supplementary Material, we show the premultiplied vertical velocity spectra ($S_{ww}(f), f=f$= frequency) from the sonic anemometer observations at locations just above the forest, at horizontal distances of one and two canopy heights ($1h$ and $2h$, where $h=$ canopy height) from the forest edge, and at the clearcut. Except the above forest one, the spectra from the rest of these locations behave very similarly for the along-wind cases (Fig. S3c) and for the cases when the wind approached the forest edge from the clearcut (Fig. S3b). In particular, under the surface layer scaling, the spectra at the clearcut and at distances one and two canopy heights collapsed reasonably well. This is not true for the wind direction from the forest edge, where the spectra at all these locations showed a poor collapse under the same scaling and remained quite different from one another (Fig. S3a). Upon a closer look at Fig. S3a, corresponding to lower frequencies, $fS_{ww}(f)$ values at a distance of $1h$ were comparable to the above forest ones, and gradually decreased with increasing distances from the forest edge.
3.1.2 Influence of sweeping effects
From Figs. 4b–c, one important aspect to consider is that both the along- and perpendicular-wind scenarios are highly turbulent, with $\sigma_{u}/U$ values exceeding 0.5. flesch_wind_1999 observed highly turbulent wind conditions in a forest clearcut approximately at a distance of three to five times the canopy height. This distance matches with the location of our EC set up, situated approximately five times the canopy height from the forest edge. In such highly turbulent conditions, for along-wind cases and when the direction of the mean wind is from the clearcut location, $U_{C}$ values, corresponding to the range $0.5<\sigma_{u}/U<0.8$, increase with $\sigma_{u}/U$. everard2021sweeping pointed out that in highly turbulent conditions, the estimations of $U_{C}$ are influenced by sweeping effects associated with large eddies of the flow. They postulated that the occurrences of the sweeping events increase considerably when the TKE of the flow is quite high. To obtain an estimate of the convective speeds, uncontaminated by the sweeping effects, they suggested to use $U_{e}$ instead (see Fig. 3).
For our clearcut conditions, in Fig. 4d, we show a scatter plot between $U_{e}$ and $U$. One immediately recognizes two distinct features. First, contrary to Fig. 4a, in Fig. 4d, a majority of data points lie below the 1:1 line, thereby suggesting that the $U_{e}$ values remain mostly smaller than both $U$ and $U_{C}$. This is exactly the same conclusion reached by everard2021sweeping for their flow configuration over a vineyard canopy. Second, the scatter among the points in Fig. 4d remains considerably larger than in Fig. 4a. This is a similar feature observed by everard2021sweeping for their vineyard canopy dataset, where they found the relationship between $U_{e}$ and $U$ did not exactly follow the 1:1 line (see their Fig. 3a). However, within the limit of experimental accuracy, they ignored this scatter and considered $U_{e}\approx U$. Specific to our set up, for the along-wind conditions (Sections 1 to 5), after ensemble averaging the results in Fig. 4d, one could find that $U_{e}$ remained approximately equal to $U$. On the other hand, for the perpendicular-wind cases (shown by red squares), the number of data points in Fig. 4d is only four. This is because, nearly for all the perpendicular-wind periods, the $r_{p}$-$\tau$ relationships could not described as a best-fit straight line with enough statistical accuracy. The statistical accuracy is quantified in terms of whether the $R^{2}$ values of the linear fits were greater than 0.9. Since the computation of $U_{e}$ is tied to the elliptic approximation model of he2010small, we hypothesize that the accuracy of this model to describe the space-time correlation curves is impacted in situations when the presence of an upstream forest canopy disturbs the flow over a clearcut location. This limitation forced us to compute the sweeping speeds ($V$) for the along-wind conditions only.
everard2021sweeping showed that the square of the sweeping speeds ($V^{2}$) was intimately connected to the TKE of the flow. They found this relationship to be $V^{2}\approx\rm 2.16\ TKE$ for heights above a vineyard canopy and $V^{2}\approx\rm 2.25\ TKE$ for heights within the canopy (see the red and magenta dash-dotted lines in Fig. 4e). To test their prediction, in Fig. 4e, we show the scatter plot between $V^{2}$ and TKE for along-wind cases. Note that the TKE is independently estimated from the EC system at the middle of a clearcut, defined as $(\sigma_{u}^{2}+\sigma_{v}^{2}+\sigma_{w}^{2})/2$. Similar to everard2021sweeping, we do see a strong relationship between the two. However, as opposed to everard2021sweeping, the relationship in Fig. 4e emerges to be $V^{2}\approx\rm TKE$ (black dash-dotted line in ig. 4e). This discrepancy is consistent with the expectation that the boundary conditions on the flow could potentially alter the relationship between $V^{2}$ and TKE (everard2021sweeping). Notwithstanding this difference, in agreement with everard2021sweeping, one can conclude that in highly turbulent conditions, the sweeping effects indeed become quite important and cause the $U_{C}$ values to appear significantly larger than the mean wind speed. The sweeping effects are also associated with modifying the turbulence frequency spectrum (everard2021sweeping), which will become apparent when we discuss the temporal to spatial conversion of turbulence spectra using scale-dependent convection speeds (see Sect. 3.2).
@@ 3.1.3 Scale-dependent convection speeds
Figure 5: (a) The scale-dependent convection speeds ($U_{C}(\kappa z)$) are plotted against the streamwise wavenumbers ($\kappa$) computed directly from the DTS spatial information. The wavenumbers are normalized by the measurement height ($\kappa z$) while $U_{C}(\kappa z)$ values are scaled by the convection speed $U_{e}$. The gray- and red-colored lines denote the along- and perpendicular-wind cases, respectively. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. (b) Same as in (a), but the $U_{C}(\kappa z)$ values are scaled by $U_{C}$. The various fits to the scale-dependent convection speeds are shown by the dash-dotted lines (see the legend).
@@
Hitherto, we have discussed the bulk convective speeds but now we turn our attention towards the scale-dependent convection speeds. In Fig. 5a, we show the ensemble-averaged scale-dependent convection speed estimations from the spatial DTS data of temperature at all the six sections. As before, the Sections 1–5 correspond to along-wind cases while Section 6 represents the perpendicular ones. The convection speeds in Fig. 5a are normalized by the uncontaminated form of the bulk convective speed, $U_{e}$, and the wavenumbers ($\kappa$) computed from the spatial DTS separations are scaled with the measurement height ($z$). As $U_{e}$ values have significant run-to-run variations (Fig. 4d), we used an ensemble-averaged value of $U_{e}$ in Fig. 5a, separately for the six different sections. Specific to the perpendicular-wind section, this average was carried out only over four 30-minute periods (see the red squares in Fig. 4d). The choice of $z$ to scale $\kappa$ is in line with hilland2024systematic, who claimed that in the context of Taylor’s hypothesis the measurement height is indeed an important parameter. Ideally, one should use zdz-d, where $d$ is the displacement height, but here we used only $z$ since to estimate $d$ a detailed vertical profile of wind measurements is required and that was not available at our measurement site. This would not have any impact on the results as $d$ only acts to shift the $x$ coordinates horizontally.
@@
Figure 5a depicts that at lower wavenumbers, $\kappa z<1$, $U_{C}(\kappa z)$ values tend towards nearly two to three times the values of $U_{e}$, irrespective of the wind directions. On the other hand, if one uses $U_{C}$ to scale the $U_{C}(\kappa z)$ values, at lower wavenumbers, $U_{C}(\kappa z)\to U_{C}$ (see Fig. 5b). At present we do not understand this difference and hence continue with normalizing the $U_{C}(\kappa z)$ values with $U_{C}$, despite acknowledging its limitations. At larger wavenumbers, $\kappa z>1$, $U_{C}(\kappa z)/U_{C}$ values decrease in a power-law fashion with the exponent being equal to ${(\kappa z)}^{-0.8}$ (Fig. 5b). This is surprisingly different from the result of cheng2017failure who reported this exponent to be ${(\kappa z)}^{-1/3}$. We replotted the results of cheng2017failure and indeed for different atmospheric stability conditions (near-neutral to highly-convective) their data agree more closely with ${(\kappa z)}^{-1/3}$ instead of ${(\kappa z)}^{-0.8}$ (see Fig. S4 in the Supplementary material). However, compared to ours, the data of cheng2017failure exhibited a significant amount of scatter. Before we discuss potential reasons behind the disagreement with cheng2017failure, it is important to recognize that, like any other observations, the DTS measurements are not perfect and are to a certain degree contaminated by noise and possibly also other instrument artifacts (peltola2021suitability). In Appendix B, we consider the impact of such noises on the 0.8-0.8 scaling exponent.
The apparent disagreement with the results of cheng2017failure could be attributed to two factors. First, the presence of anisotropy at smaller scales of the flow. cheng2017failure theoretically derived their 1/3-1/3 exponent for atmospheric surface layer flows by using the advection-diffusion equation and a power-law dependence in the turbulent diffusion coefficient at the inertial subrange scales. In fact, they considered the turbulent diffusion coefficient of the inertial subrange eddies to vary as $r^{4/3}$, where $r$ represent their sizes. This $4/3$ power-law is a direct consequence of the turbulent particle pair diffusion theory introduced by Richardson (richardson1926atmospheric). richardson1926atmospheric assumed that the turbulence is locally isotropic at the inertial subrange scales, and the energy spectrum follows a 5/3-5/3 slope. In recent times, malik2018turbulent has shown that in the limit of asymptotically infinite inertial subrange (symptomatic of atmospheric flows with very-high Reynolds number), the exponent of the power-law ($\gamma$) associated with turbulent diffusion can vary between $(1+p)/2<\gamma<2$, where $p$ is the slope of the energy spectrum in the inertial subrange. If $p$ is $5/3$, then this range turns out to be between $4/3<\gamma<2$.
While deriving this result, malik2018turbulent questions the assumption of locality by introducing a hypothesis that the non-local scales of motion do contribute to the particle pair diffusion process at inertial subrange scales. In the definition of malik2018turbulent, the sizes of these non-local eddies are larger than the inertial subrange ones but smaller than the largest scales, typically the integral scales of streamwise velocity fluctuations. Now, in our case, we obtain a power-law of 0.8-0.8. If one follows the same line of derivation as outlined in cheng2017failure, a 0.8-0.8 power-law can be obtained if the turbulent diffusion coefficient of the inertial subrange eddies could be expressed as, $r^{1.8}$. Although this range of $\gamma$ can justify a relationship of $r^{1.8}$ (since $4/3<1.8<2$), from point-based temporal measurements in atmospheric flows, it is nearly impossible to provide direct evidence of this non-locality. However, this can be done indirectly. The presence of non-locality at the inertial subrange scales impart anisotropy at those scales (katul1997energy). The results of chowdhuri2024quantifying have demonstrated that, for a wide range of both atmospheric and laboratory datasets, significant anisotropy can persist even at smaller scales of the flow. This is true for our EC dataset, as the flow remains anisotropic at the inertial subrange scales when computed through the intermittency-anisotropy framework of chowdhuri2024quantifying (not shown).
The second reason is related to turbulence intensities. For our heterogeneous flow set up (a forest clearcut), the conditions are highly turbulent with $\sigma_{u}/U>0.5$. cheng2017failure did not report the values of their turbulence intensities. Later on, han2022applicability reevaluated the dataset of cheng2017failure. They reported that the $\sigma_{u}/U$ values of the eight half-hour runs of cheng2017failure (spanning different stability conditions) were less than 0.5 with only one exception (see Table 1 of han2022applicability). Thus, cheng2017failure mostly encountered low turbulent wind conditions ($\sigma_{u}/U<0.5$). The study by everard2021sweeping has shown that in highly turbulent conditions, significant sweeping effects associated with large eddies can persist in the flow and can induce incorrect spectral scalings when converting the spectra from frequency to wavenumber domain by using only the mean wind speed. This indeed can be accounted for by using scale-dependent convection speeds. everard2021sweeping established this by considering a limiting case with zero-mean wind speed but with non-zero sweeping effects (see their Eqs. 6 and 9). Therefore, apart from anisotropic effects at smaller scales, the bulk turbulence intensities could also have its impact on the exponents of scale-dependent convection speeds. Next, we provide a recipe to convert the spectra from temporal to spatial domain by using scale-dependent convection speeds.
3.2 Conversion of turbulence spectra